What are the odds of two people flopping a straight flush poker?

  • Thread starter Thread starter PokerSki
  • Start date Start date
I was confused ,as if you meant in the same hand or just at the same table in different times of the game ,lol ,,,, in the same hand and both on the flop? I honestly don't think that would ever happen the odds are probably astronomical you probably have more chance of winning a huge Powerball lottery🙂
 
The odds of being dealt a straight flsuh are 1 in 64,974 as for being dealt one on the flop.. Not sure.. It depends on your hole cards. The best odds are with mid range suited connectors, specifically 45s to JTs. For example, with 45s, you have 4 possible ways of flopping a straight flush (A23s, 236s, 367s, and 678s) for both I don't know but I saw in a video of a wsop hand one guy flopped a royal flush and his opponet rivered 4 aces and when he hit the rail the announcer said it was 3.1 billion to 1...but it happened. It's on you tube.
 
So just assuming the basic odds of hitting a straight flush, the chance of two players having one is .00000000024%. That number is even smaller taking into account that both players have to be dealt specific suited connectors such that they both can have a straight flush at the same time, and delimiting it to just the flop gets us into the trillions to one level. In all my years of playing poker I have been involved in one straight flush over straight flush hand in Holdem and two in Seven Stud.
 
flop

2 on flop 0 not possaible .havefunAA ok went back and looked you said str f iwas thinging rf so could heappen but astromakiel odds still real slim .havefunAA
 
Last edited:
I had this exact scenario happen to me! I had 9h10h, he had 4h5h, flop came 678 all hearts. I'm a poker dealer and this is still to date the craziest hand I've ever seen and will likely ever see. Anyone able to compute the odds of this happening?! I'm all too curious
 
  • Like
Reactions: tcummo
The odds are very small for getting such combinations, I once had such a situation in the game that I caught a straight flush and another player also had a down flash straight flush combination I had 67 tambourines on hands and the had 8- on the table 9-10 tambourine - and the player had JQ tambourine cards and we got two straight flush but I lost the hand!
 
I think no one and never flopping a straight flush in poker is a very strong combination, and the probability that both will have a straight flush is very small.
 
it is not possible to give an exact answer without knowing the pocket card.
for exampleA2 or AK suited hasless chance than 78 suited of hitting the st flush
 
As often happens I spent an hour replying to a thread only to time out before I can post it. Alas you are all rewarded with the cliff notes version of the original.

Odds of you specifically losing with a straight flush to a royal flush are 77260583400:1,
but the AK vs 98 situation is only 1 of 8 SF over SF situations so the odds drop to 9657572925:1.

But your question wasn't about you specifically losing SF over SF but rather witnessing it at the table. Assuming that there are ten players at the table and each player has an equal chance of drawing the short end of the straw I will say that the odds increase by a factor of 20 bringing the likelihood of witnessing such an event down to a rather mundane 482878646.25:1.

Playing live at a zippy 25 hands per hour and say you play 10 hours/week you should see this about one every 38000 years, so the last time it happened would have been at that awesome home game in your cave at Lascaux.
 
  • Like
Reactions: dannystanks
Yesterday I made a straight flush on the turn, my opponent made a straight flush older than mine. When I called him allin I thought what chances he would have a straight flush too. It turned out that the odds are 100%​
 
If you play at Pokerstars (or any other client that involves money) you must have imagined that 92.86% probability of winning is just a mathematical number so that we can better tolerate (or quantum mechanics) the variances.

Hand

:kc: :flute:
 
DTM22 said:
I had this exact scenario happen to me! I had 9h10h, he had 4h5h, flop came 678 all hearts. I'm a poker dealer and this is still to date the craziest hand I've ever seen and will likely ever see. Anyone able to compute the odds of this happening?! I'm all too curious
To add the math, (but it depends the interpretation of the question):
the flop is given: 3 card in a row of same color, with lowest >3 and highest < K
player A has the 2 lower hole cards, chance: 1/49 * 1/48 = 1/2352
player B has the 2 higher hole cards, chance: also 1/2352
The chance of both happening in the same hand: 1/(2352*2352) = 1/5531904
Note:
if you flop the lower straight flush like this: the chance of someone else having the better straight flush is 1/2352
 
I had 4d5d pocket, flop came 3d6d7d, two players went all-in and I called, first player shows AdJd, second player shows 8d9d, the turn was 10d. Not all on the flop, but that is the closest I seen.
 
dedok0525 said:
Yesterday I made a straight flush on the turn, my opponent made a straight flush older than mine. When I called him allin I thought what chances he would have a straight flush too. It turned out that the odds are 100%​
According to Murphy's law, that's correct. Better fold it next time!…………………………… J/k <; according to the same law, a fold there surely means that you folded the best hand.

Seriously, considering how more than half of the 5 cards of each player are shared on the flop by all players in a poker game like Texas Holdem (involving community cards), it's more likely that the players will hold similar 5 card hands (flushes, straights), so flush vs flush and straight vs straight should be more common than one would think.
 
8bod8 said:
To add the math, (but it depends the interpretation of the question):
the flop is given: 3 card in a row of same color, with lowest >3 and highest < K
player A has the 2 lower hole cards, chance: 1/49 * 1/48 = 1/2352
player B has the 2 higher hole cards, chance: also 1/2352
The chance of both happening in the same hand: 1/(2352*2352) = 1/5531904
Note:
if you flop the lower straight flush like this: the chance of someone else having the better straight flush is 1/2352
The above numbers are 'not correct', looking a bit closer. Corrected, for heads up:
player A has the 2 lower hole cards, chance: 4/49 * 1/48 = 1/588
player B has the 2 higher hole cards, chance: 2/47 * 1/46 = 1/1081
The chance of both happening in the same hand: 1/(588*1081) = 1/635628
Note, not only for heads up:
- if you flop the lower straight flush like this: the chance of someone else having the better straight flush is 1/1081, or 0.1% per other player in the hand.
- if the flop shows as above, and you don't block any straight flush, the chance of each player still in the hand having a straight flush is 1/588, or 0.17%. example: with a 7 way flop, there is 6/588 = 1% chance of someone else having a straight flush.
 
I had and instance where I had 8d9d runout had 10d Jd Qd vs AdKd. I had turned the straight flush and river brought the royal to the villain.
This was at a small tournament at a bar in Prescott, Az. I had brought up I suspected some shady stuff happening over the period of a hand full of tournaments I entered. I guess this was their way of letting me know I wasn’t welcomed. To be honest when I turned the SF, I already knew what was going to happen on the river.
 
The odds are approximately 1 in 439,000,000.
Calculations could possibly be off, but they seem accurate.

Love me some math ............
 
I remember a game many years ago from UK tv show Late night poker with 3 maybe 4 players in the pot and the table came with a straight flush (I think it was a king high) one of the players had the A. So more than 2 player have had straight flushes in the same hand. and lost.
 
Back
Top